### v1

# CMRAG: Co-modality–based document retrieval and visual question answering

Abstract
Retrieval-Augmented Generation (RAG) has become a core paradigm in document question answering tasks. However, existing methods have limitations when dealing with multimodal documents: one category of methods relies on layout analysis and text extraction, which can only utilize explicit text information and struggle to capture images or unstructured content; the other category treats document segmentation as visual input and directly passes it to visual language models (VLMs) for processing, yet it ignores the semantic advantages of text, leading to suboptimal generation results. This paper proposes co-modality–based RAG (CMRAG), which can simultaneously leverage text and images for efficient retrieval and generation. Specifically, we first perform structured parsing on documents to obtain co-modality representations of text segments and image regions. Subsequently, in response to user queries, we retrieve candidate evidence from text and image channels, respectively, and aggregate the results at the cross-modal retrieval level. Finally, we prompt the VLM to generate the final response based on the co-modality retrieval results. Experiments demonstrate that our method significantly outperforms pure-vision–based RAG in visual document question answering tasks. The findings of this paper show that integrating co-modality information into the RAG framework in a unified manner is an effective approach to improving the performance of complex document visual question-answering (VQA) systems.
1 Introduction
Figure 1: Comparison among (a) text–based RAG, (b) image–based RAG, and (c) co-modality–based RAG.
Large language models (LLMs) have received extensive attention in recent years (Touvron et al., 2023; Achiam et al., 2023; Guo et al., 2025; Yang et al., 2025a), but they have inherent limitations in handling out-of-domain knowledge (Ji et al., 2023). To address this issue, RAG integrates external knowledge retrieval with the generation process (Lewis et al., 2020; Guu et al., 2020; Karpukhin et al., 2020; Chen et al., 2025). RAG achieves wide success in open-domain question answering, knowledge retrieval, and dialogue systems, and becomes an effective means of extending the knowledge boundaries of LLMs (Ram et al., 2023; Gao et al., 2023). However, most external data sources (e.g., documents) are essentially multimodal (Jeong et al., 2025; Faysse et al., 2025), often containing natural language text, formulas, tables, images, and complex layout structures. How to effectively leverage such multimodal information in question answering remains a challenging problem that is not fully solved.
One line of approaches is text-based RAG, which typically relies on layout parsing and text extraction (Xu et al., 2020; Dong et al., 2025; Yang et al., 2025b; Perez & Vizcaino, 2024). These methods first detect document layouts and then extract textual information for subsequent retrieval and generation. While stable at the semantic level, they struggle to handle content such as images and tables. Recently, VLMs (Ghosh et al., 2024; Radford et al., 2021; Alayrac et al., 2022) enable RAG systems to process documents directly as images (Faysse et al., 2025; Yu et al., 2025; Qi et al., 2024a; Wang et al., 2025b), giving rise to vision-based RAG (Huang et al., 2022; Kim et al., 2022; Yu et al., 2024). Specifically, these methods divide document pages into image segments and perform retrieval and reasoning through visual understanding models. Although they capture non-textual information, they often overlook the precise information carried by text, leading to performance bottlenecks.
To overcome these limitations, we propose a novel co-modality–based RAG framework, which unifies text and image modalities, as illustrated in Fig. 1. In this framework, we first parse documents to extract structured text and image segments. For a given query, we then perform retrieval in both text and visual spaces, ensuring that semantic matching from text and perceptual grounding from images are simultaneously leveraged. Finally, we feed the co-modality evidence into a VLM to integrate information and generate answers. In this way, our method not only handles cross-modal reasoning that single-modal RAG cannot cover but also achieves significant performance improvements on standard visual document question answering benchmarks.
The main contributions of this paper are summarized as follows:
  • We propose a novel RAG framework that jointly leverages text and image modalities to enhance performance on complex visual document QA tasks.
  • We design a co-modality–based retrieval mechanism that enables complementary use of text and image evidence during retrieval, leading to more comprehensive knowledge grounding.
  • We conduct extensive experiments on document VQA benchmarks, and the results demonstrate that our approach significantly outperforms existing single-modal RAG methods in both accuracy and robustness.
2 Related work
In this section, we discuss the recent studies related to multi-modal RAG (MMRAG). Specifically, we primarily focus on three folds: (1) knowledge–based MMRAG, (2) video–based MMRAG, and (3) document–based MMRAG.
Knowledge–based MMRAG refers to retrieving knowledge (text or image modality) from external sources such as Wikipedia articles and websites to answer textual or visual queries (Talmor et al., 2021; Marino et al., 2019; Chang et al., 2022; Schwenk et al., 2022; Mensink et al., 2023; Chen et al., 2023; Ma et al., 2024a; Hu et al., 2025). Although the external knowledge database can enhance the system performance (Caffagni et al., 2024), the key issue of knowledge–based MMRAG is the inconsistency between textual and visual queries as well as the external knowledge database (Chen et al., 2022; Lin et al., 2023; Zhang et al., 2024b). To address this issue, Lin & Byrne (2022) adopted multiple algorithms, including object detection, image captioning, and OCR, to transform visual queries into language space, and proposed a joint training scheme to optimize retrieval and generation simultaneously. A similar training strategy was also used by (Adjali et al., 2024). Also, Yan & Xie (2024) used a consistent modality for both retrieval and generation: visual modality for retrieval (visual queries and Wikipedia article pages) and textual modality for generation (textual queries and wiki articles). A similar strategy can also be found in RORA (Qi et al., 2024a). In addition, Tian et al. (2025) proposed cross-source knowledge reconciliation for MMRAG, which could address the inconsistency between textual and visual external knowledge.
Video–based MMRAG refers to retrieving videos from the corpus to help answer given queries (Caba Heilbron et al., 2015; Xu et al., 2016; Anne Hendricks et al., 2017; Wang et al., 2019; Kriz et al., 2025; Wan et al., 2025). Since encoding videos may incur high computational costs, a few studies pre-processed videos using VLMs and converted videos to textual modality (Zhang et al., 2024a; Arefeen et al., 2024; Ma et al., 2025). For example, Zhang et al. (2024a) first detected key information in videos such as human faces, based on which a VLM was prompted to generate scene captions for video frames. Consequently, the video modality can be converted to a text modality, which can significantly reduce computational costs and facilitate retrieval and generation. Furthermore, a few studies (Luo et al., 2024; Jeong et al., 2025; Reddy et al., 2025) processed videos by selecting or clustering representative frames, facilitating video retrieval and final generation.
Document–based MMRAG refers to retrieving a few document pages from one or multiple documents to help generate answers for given questions (Methani et al., 2020; Masry et al., 2022; Tanaka et al., 2023; Tito et al., 2023; Ma et al., 2024b; Li et al., 2024; Hui et al., 2024; Qi et al., 2024b; Cho et al., 2024; Wang et al., 2025a; Li et al., 2025; Wasserman et al., 2025; Faysse et al., 2025). Traditionally, documents were parsed using detection models (Ge et al., 2021) and OCR engines (Smith, 2007), and the extracted components (e.g., text) were input to LLMs to generate answers (Riedler & Langer, 2024; Faysse et al., 2025). With the proliferation of VLMs, a few studies (Faysse et al., 2025; Yu et al., 2025; Wang et al., 2025b) processed document pages as images directly. Specifically, they used VLMs to encode queries and document pages as text and images, respectively, based on which the similarity scores between queries and visual document pages can be calculated. This method paves the way for document–based MMRAG, as it does not need to parse documents and can retrieve document pages directly. However, this method overlooks text modality in documents, which may degrade the performance of RAG systems.
3 Methodology
In this section, we introduce the problem definition of document–based VQA and our proposed CMRAG framework in detail.
3.1 Problem definition
In a document–based VQA system, a collection of visual documents provides evidence for answering given queries. Formally, all pages are treated as candidate evidence $\mathcal{D}=\{p_{1},p_{2},\dots,p_{M}\}$, where $M$ denotes the total number of pages across the document collection. For a given query $q$, the retriever $\mathcal{R}$ identifies the top-$k$ relevant pages $P_{k}=\{p_{1},p_{2},\dots,p_{k}\}$ according to a similarity score $s(q,p_{i})$ ($\forall p_{i}\in P_{k}$). Once the top-$k$ evidence pages are retrieved, they are combined with the query into a prompt $\mathcal{P}(q,P_{k})$, which is then passed to a generator model $\mathcal{G}$, often instantiated as a vision–language model (VLM). The generator integrates the multimodal evidence and produces the final answer $\hat{a}=\mathcal{G}(\mathcal{P}(q,P_{k}))$.
3.2 VLM–based document parsing
Advanced vision–language models (VLMs) can directly parse document pages, simultaneously detecting layouts, extracting text, and identifying localized visual regions. For each page $p_{i}$, the parsing process produces three complementary components: the entire image $I_{i}$, which preserves global layout and contextual relationships; the parsed sub-images $SI_{i}$, which capture localized elements such as figures, tables, and diagrams; and the parsed text $T_{i}$, which provides precise semantic grounding for fine-grained reasoning. Thus, each page can be represented as $p_{i}=\{I_{i},SI_{i},T_{i}\}$, ensuring that both global structure and detailed content are available for retrieval and downstream answer generation.
3.3 Co-modality late interaction
Late interaction. Given a query $q$ and document page image $p$, the encoder $\mathcal{E}$ project them into a high-dimension space $\mathbb{R}^{D}$, i.e., $\mathbf{q}=\mathcal{E}(q)\in\mathbb{R}^{N_{q}\times D}$ and $\mathbf{p}=\mathcal{E}(p)\in\mathbb{R}^{N_{p}\times D}$, where $N_{q}$ and $N_{p}$ denote the number of vectors in the query and document embeddings, respectively. Hence, the late interaction (LI) (Khattab & Zaharia, 2020; Santhanam et al., 2022; Clavié, 2025; Faysse et al., 2025) can be calculated as follows:
$\textbf{LI}(q,p)=\sum_{i=1}^{N_{q}}\max_{j\in\{1,2,...,N_{p}\}}\langle\mathbf{q}^{i},\mathbf{p}^{j}\rangle,$ (1)
where $\langle\cdot,\cdot\rangle$ denote dot production, and $\mathbf{q}^{i}$ and $\mathbf{p}^{j}$ are $i^{\text{th}}$ and $j^{\text{th}}$ vector in the query and page embeddings, respectively. Compared with the mean pooling method (Karpukhin et al., 2020; Yu et al., 2025), LI can consider token-level similarity for detailed matching (Wan et al., 2025).
Co-modality late interaction. Furthermore, inspired by CLAMR (Wan et al., 2025), we consider retrieving evidence based on similarity between the query $q$ and the co-modality document content $I$ and $T$. Formally, we encode the entire page image $I$ and extracted text $T$ as $\mathbf{p}^{I}=\mathcal{E}(I)$ and $\mathbf{p}^{T}=\mathcal{E}(T)$, respectively. Therefore, we adopt co-modality late interaction (CoLI) to calculate the similarity, as follows:
$\textbf{CoLI}(q,p)=\max_{m\in\{I,T\}}\sum_{i=1}^{N_{q}}\max_{j\in\{1,2,...,N_{p}\}}\langle\mathbf{q}^{i},\mathbf{p}^{m,j}\rangle,$ (2)
where $\mathbf{p}^{m,j}$ is $j^{\text{th}}$ vector in the $m^{\text{th}}$ modality document embedding.
Contrastive learning. Let $\{(q_{i},p_{i})\}_{i=1}^{b}$ denote a data batch where each query $q_{i}$ ($\forall i\in\{1,2,...,b\}$) corresponds to a document page $p_{i}$ and $b$ denotes the batch size. The loss (Oord et al., 2018) can be represented as follows:
$\mathcal{L}=-\frac{1}{b}\sum_{i=1}^{b}\log\frac{\exp(s_{i,i}/\tau)}{\sum_{j=1}^{b}\exp(s_{i,j}/\tau)},$ (3)
where $s_{i,j}$ ($\forall i,j\in\{1,2,...,b\}$) denote the similarity score between $i^{\text{th}}$ query and $j^{\text{th}}$ document page and $\tau$ the temperature.
3.4 Co-modality–based generation
Recent advances in VLMs (Bai et al., 2025; Du et al., 2025) enable them to directly process naive-resolution page images without the need for splitting each page into smaller sub-images, which often risks losing global context. In addition, modern VLMs can accept multiple images simultaneously, meaning that all retrieved pages can be jointly fed into the model. This design facilitates cross-page reasoning, where evidence from different parts of the document can be integrated to produce more accurate and comprehensive answers. Therefore, we combine the retrieved images and text into the prompt and leverage the VLM to generate the final answer based on the co-modality input.
4 Experiments
4.1 Experiment setup
Datasets. We consider two datasets: MMLongBench (Ma et al., 2024b) and SlideVQA (Tanaka et al., 2023). These datasets comprise various query types (multi-hop and single-hop) and answer types (multi-span and single-span), which can be used to evaluate the capacities of VLMs. The statistics of the datasets are summarized in Tab. 1. We follow the implementation of (Wang et al., 2025a) and use refined queries of SlideVQA.
Method #Files #Images #QAs #Images/Q
MMLongBench 135 6,529 1,082 1.00
SlideVQA 400 8,000 2,020 1.26
Table 1: Statistics of evaluation datasets.
Baselines. To evaluate the proposed method, we first consider three evaluation scenarios: (1) Oracle (Entire images), (2) Oracle (Sub-images + parsed text), and (3) Oracle (Entire images + parsed text). We consider using the ground truth document pages, with the entire images, parsed sub-images plus text, and entire images plus parsed text as the input of the VLM, respectively.
LLM-as-a-Judge. The gold answers in the benchmark datasets consist of multiple spans or exhibit diverse valid expressions. Traditional evaluation metrics, such as Exact Match (EM) and token-level F1, struggle to capture the semantic equivalence between generated responses and these multi-span gold answers. As a result, these metrics may underestimate the actual performance of models. To address this limitation, we employ a large language model (LLM) as an automatic judge to assess whether a generated answer is correct, providing a more flexible and reliable evaluation framework, enabling fairer comparisons across different models.
Implementation details. We implemented our CMRAG framework using Qwen2.5-VL-7B-Instruct Bai et al. (2025) as the backbone VLM for parsing documents and answering queries. In addition, we adopted Qwen2.5-7B-Instruct (Yang et al., 2025a) as the judge. To ensure deterministic outputs and eliminate variability due to random sampling, we set the temperature to 0.0 during decoding. We present the prompts used in this study in the appendix.
Method MMLongBench SlideVQA
Oracle (Entire images) 40.39 81.58
Oracle (Sub-images + parsed text) 29.11 58.76
Oracle (Entire images + parsed text) 46.86 81.68
Table 2: Main results. Bold values represent the best scores.
4.2 Main results
Tab. 2 summarizes the performance of different oracle settings on MMLongBench and SlideVQA. The results reveal several key findings. First, incorporating parsed text consistently enhances generation accuracy, as shown by the improved scores of the oracle that integrates entire images with parsed text. For instance, on MMLongBench, the combined setting achieves the best score of 46.86, which significantly outperforms using only entire images (40.39). Similarly, on SlideVQA, the combination yields 81.68, slightly higher than the image-only oracle. Second, directly using sub-images with parsed text leads to degraded performance. This decline is likely due to incomplete visual information within sub-images, which may omit contextual elements crucial for accurate reasoning. On MMLongBench, the score drops to 29.11, while SlideVQA falls to 58.76—both substantially lower than their entire image counterparts. These results suggest that incomplete visual segmentation can hinder the integration of multimodal cues. Overall, the findings highlight that while parsed text provides a complementary signal that improves multimodal reasoning, extracting sub-images without preserving global context can be detrimental. The best results are achieved when both entire images and parsed text are jointly considered.
Figure 2: Qualitative comparison among three baselines.
4.3 Case study
To further illustrate the effects observed in Table 2, we present a case study in Fig. 2. This example highlights the strengths and limitations of different Oracle settings in handling complex queries. First, when relying solely on entire images, the VLM misinterprets the numbers, producing an incorrect output of 36 instead of 62. This demonstrates that VLMs may struggle to accurately ground numeric reasoning based on purely visual inputs. Similarly, when only sub-images plus parsed text are used, the model fails to capture the complete context, yielding partial and incomplete answers. The problem arises because the VLM failed to accurately parse the sub-image but extracted textual numbers only. However, incorporating entire images together with parsed text enables the model to generate the correct multi-span answer, as the parsed text provides a reliable textual grounding that compensates for the VLM’s difficulty in interpreting fine-grained visual details. This shows that parsed text can serve as an essential complement, ensuring accurate reasoning across multiple evidence spans.
Another challenge revealed by this case is that the query requires multi-span answers, i.e., identifying two distinct values from different textual locations. Conventional automatic evaluation metrics, such as exact match (EM) or F1 score, cannot adequately capture the correctness of such answers, which further validates the necessity of LLM-as-a-Judge.
5 Conclusion
In this paper, we introduced a co-modality–based RAG framework that unifies text and image modalities to address the limitations of existing single-modal approaches in visual document question answering. By jointly leveraging parsed text and entire images, our method enables complementary retrieval and reasoning, where textual evidence provides precise grounding and visual context preserves global completeness. Experimental results on MMLongBench and SlideVQA demonstrate that the proposed framework consistently outperforms single-modality baselines, and case studies further highlight its advantages in handling multi-span queries that traditional metrics cannot fully capture. Overall, this work establishes co-modality RAG as an effective and robust paradigm for document VQA, paving the way for future research on integrating structured parsing, fine-grained retrieval, and LLM-based evaluation.
Appendix A Prompt template
Figure 3: Prompt templates for (a) parsing images, (b) generating answers based on entire images, (c) generating answers based on sub-images and text, generating answers based on entire images and text, and (e) judging generated answers. The first template can be found at https://github.com/QwenLM/Qwen2.5VL/blob/main/cookbooks/document_parsing.ipynb and the rest can be referred to (Wang et al., 2025b).